微分幾何学 III （大学院） 講義ノート

参考文献 [岡]岡本清郷, 等質空間上の解析学, 紀伊國屋書店, 東京, 1980. [竹]竹内勝, 現代の球関数, 岩波書店, 東京, 1975. [松]松島与三, 多様体入門, 裳華房, 東京, 1965. [横]横田一郎, 群と表現, 裳華房, 東京, 1973.

[A] J. F. Adams, Lectures on Lie groups, Univ. Chicago Press, Chicago, 1969.

[B] N. Bourbaki, Groupes et algèbres de Lie, Hermann, Paris, 1975. [B-D] T. Bröcker and T. t. Dieck, Representations of compact Lie groups, Springer,

New York, 1985. [C-E] J. Cheeger and D. G. Ebin, Comparison theorems in Riemannian geometry,

North-Holland, Amsterdam, 1975.

[D] J. Dieudonné, Foundations of modern analysis (Enlarged and corrected printing), Academic Press, New York, 1969.

[F] H. Federer, Geometric measure theory, Springer, Berlin, 1969.

[H] S. Helgason, Groups and geometric analysis, Academic Press, New York, 1984. [M]J. W. Milnor, Topology from the differentiable viewpoint, University Press of

Virginia, 1965. [W]N. R. Wallach, Harmonic analysis on homogeneous spaces, Marcel Dekker, New

York, 1973.